OpenStudy (anonymous):

Solve the equations: x^2+y^2=13, xy+6=0

OpenStudy (anonymous):

using eqn 1 x=-6/y then substitute in eqn 2 (-6/y)^2+y^2=13

OpenStudy (anonymous):

whats after this ??

OpenStudy (callisto):

Multiply both sides by y^2

OpenStudy (callisto):

So, you'll get: \[(\frac{-6}{y})^2+y^2=13\]\[(-6)^2+y^4=13y^2\]Quadratic equation..

OpenStudy (anonymous):

it has y^4 how is this quadratic?

OpenStudy (callisto):

\[(y^2)^2 - 13y^2 + 36=0\]

OpenStudy (callisto):

Let u = y^2 u^2 - 13u + 36 =0 Solve u, the solve u=y^2

OpenStudy (anonymous):

Another thing you could try: \[ x=\sqrt{13}\cos t,\quad y = \sqrt{13}\sin t \\ 13\sin t\cos t+6=0 \implies \frac{13}{2}\sin 2t=-6 \]

OpenStudy (badhi):

$$(x^2+y^2)=13$$$$xy=-6\implies 2xy=-12$$ adding those two together, $$x^2+2xy+y^2=1\implies (x+y)^2=1\implies x=-y\pm1$$ considering plus, $$x=1-y\\ (1-y)y=-6\implies y^2-y-6=0\implies y=3,\,y=-2$$so, $$x=-2,\, x=3$$ consider minus, $$x=-(y+1)\\(y+1)y=6\implies y^2+y-6=0\implies y=-3,\,y=2$$ so, $$x=2,\,x=-3$$ the answers are, $$(-2,3),\,(3,-2),\,(2,-3),\,(-3,2)$$